# Optimizing Temperature Distributions for Training Neural Quantum States using Parallel Tempering

### v2

@@ Acknowledgements.
We would like to thank the UNM Center for Advanced Research Computing, supported in part by the National Science Foundation, for providing the high performance computing resources used in this work. This material is based upon work supported by the National Science Foundation under grant number 2037755. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-22-1-0498. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. This article has been co-authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan. The Flatiron Institute is a division of the Simons Foundation.
@@ Appendix C Hyperparameter dependence in the $J_{1}$-$J_{2}$ problem
Finally, we also investigate varying the number of replicas within the PT algorithm in Fig. 9. We find a strong peak in the success rate when allowing the temperatures to optimize at 6 replicas. Beyond this point the success rate generally seems to saturate near $0.60$ when the temperature is allowed to optimize. The success rate for cases where the temperature is not optimized also slowly increases to nearly match the adjusting temperature rate, demonstrating that as more temperatures are covered, the benefits of optimizing the temperature become less strong.
However, we do see more variable success rate results around 4 and 5 replicas, with 4 replicas seeing a spike in success rate when temperatures are allowed to optimize. This may be related to an optimum average swap probability, as shown in Fig. 9. As the number of replicas increases, the average swap probability increases relatively linearly for both cases where the temperature is allowed to optimize and cases where it is not. The spread of the probabilities is much narrower when the temperatures are allowed to optimize, however, confirming that the optimization algorithm is leading to a degree of convergence in the swap probability. Probabilities around $0.20$ have previously been associated with optimum success probabilities in similar PT algorithms [43]. In our work, the optimum probability of swapping appears to be around $0.40$ for the 6 replica case. This higher rate of swapping may be related to the limited number of samples and total updates performed in this work.

### v3

@@ Acknowledgements.
We would like to thank the UNM Center for Advanced Research Computing, supported in part by the National Science Foundation, for providing the high performance computing resources used in this work. This material is based upon work supported by the National Science Foundation under grant number 2037755. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-22-1-0498. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. The Flatiron Institute is a division of the Simons Foundation. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan.
@@ Appendix C Hyperparameter dependence in the $J_{1}$-$J_{2}$ problem
Finally, we also investigate varying the number of replicas within the PT algorithm in Fig. 9. We find a strong peak in the success rate when allowing the temperatures to optimize at 6 replicas. Beyond this point, the success rate generally seems to saturate near $0.60$ when the temperature is allowed to optimize. The success rate for cases where the temperature is not optimized also slowly increases to nearly match the adjusting temperature rate, demonstrating that as more temperatures are covered, the benefits of optimizing the temperature become less strong.
However, we do see more variable success rate results around 4 and 5 replicas, with 4 replicas seeing a spike in success rate when temperatures are allowed to optimize. This may be related to an optimum average swap probability, as shown in Fig. 9. As the number of replicas increases, the average swap probability increases relatively linearly for both cases where the temperature is allowed to optimize and cases where it is not. The spread of the probabilities is much narrower when the temperatures are allowed to optimize, however, confirming that the optimization algorithm is leading to a degree of convergence in the swap probability. Probabilities around $0.20$ have previously been associated with optimum success probabilities in similar PT algorithms [43]. In our work, the optimum probability of swapping appears to be around $0.40$ for the 6 replica cases. This higher rate of swapping may be related to the limited number of samples and total updates performed in this work.
@@
Appendix D Energy Comparisons
For completeness, we show in Tables 1, 2, and 3 the energies measured after $2\times 10^{4}$ training steps of the feedforward networks from Sec. III.3 for PT with temperature optimization, PT without temperature optimization, and no PT. The energies of the successful runs, defined as the subset of runs that found an energy lower than the first excited state energy, are consistent within error bars in all three cases. There is, however, a large improvement over the no PT case for the PT with and without temperature optimization cases if we consider the mean over all runs, indicating that PT on average achieves lower energies.
These results imply that the main limitation to finding a lower energy is the expressivity of the network. What PT and temperature optimization offer is the ability to more often reach the limit of the architecture’s expressivity. This is seen most clearly in Tables 2 and 3 where the total and success means for PT with temperature optimization are identical. In this case, the training procedure found a minima lower than the first excited state by the $2\times 10^{4}$ step in every simulation we performed.
Total Mean Success Mean
No-PT $-64.3\pm 0.9$ $-70.23\pm 0.03$
PT-NTO $-69.9\pm 0.1$ $-70.244\pm 0.007$
PT-TO $-70.15\pm 0.05$ $-70.253\pm 0.005$
Table 1: Mean energies for the (32,16,16) network on the $6\times 6$ $J_{2}-J_{2}$ model with $J_{2}=0.62$ over 50 independent runs. We compare no parallel tempering (No-PT) to parallel tempering without temperature optimization (PT-NTO) to parallel tempering with temperature optimization (PT-TO).
Total Mean Success Mean
No-PT $-66.4\pm 0.6$ $-70.29\pm 0.02$
PT-NTO $-70.0\pm 0.1$ $-70.314\pm 0.009$
PT-TO $-70.309\pm 0.008$ $-70.309\pm 0.008$
Table 2: Mean energies for the (48,24,24) network on the $6\times 6$ $J_{2}-J_{2}$ model with $J_{2}=0.62$ over 50 independent runs. We compare no parallel tempering (No-PT) to parallel tempering without temperature optimization (PT-NTO) to parallel tempering with temperature optimization (PT-TO).
Total Mean Success Mean
No-PT $-67.5\pm 0.4$ $-70.30\pm 0.02$
PT-NTO $-70.17\pm 0.07$ $-70.36\pm 0.01$
PT-TO $-70.36\pm 0.008$ $-70.36\pm 0.008$
Table 3: Mean energies for the (64,32,32) network on the $6\times 6$ $J_{2}-J_{2}$ model with $J_{2}=0.62$ over 50 independent runs. We compare no parallel tempering (No-PT) to parallel tempering without temperature optimization (PT-NTO) to parallel tempering with temperature optimization (PT-TO).